# Unlocking Open-Set Language Accessibility in Vision Models

### v1

@@ 4.2 Zero-Shot Decoding of Visual Features into Text
In this application, we aim to decode the visual feature vector $f$ for an image $I$, given by $f=F_{v}(I)$, into natural language sentence. This offers a text-based interpretation of what the visual features contain. We adapt the method introduced in (Tewel et al., 2021) for our purpose. Specifically, we first project the visual feature vector $f$ using the MLP to obtain $\tilde{f}$. That is, $\tilde{f}=\text{MLP}(f)$. Since $\tilde{f}$ is now in the same space as the text encoder $T$, we can measure its association to any encoded text. We utilize an off-the-self pretrained language decoder model (e.g., GPT-2), denoted as $G$, to generate open-ended text. We keep $G$ frozen to maintain its language generation capabilities and instead use prefix-tuning (Li & Liang, 2021) to guide $G$ to generate a text that maximizes the similarity with the transformed visual feature vector $\tilde{f}$. Specifically, we attach a set of learnable “virtual” tokens to $G$. Denote the generated output text of $G$ for one iteration as $h^{j}$, where $j$ represents the iteration number. $h^{j}$ is encoded with the text encoder $T$, to yield a vector $y^{j}$, That is, $y^{j}=T(h^{j})$. As $y^{j}$ and $\tilde{f}$ are now in the same embedding space, we maximize the cosine similarity between them in order to update the learnable tokens. We perform this process for several iterations. An overview of this process is shown in Figure 4. As this process is not the core contribution of our work, we leave more details to Section E of the appendix, which also includes a detailed illustration in Figure 9.
Figure 4: To decode visual features into natural language using any pretrained language decoder (e.g., GPT-2), we apply prefix tuning while keeping the language decoder fixed, generating text that maximizes the similarity with the visual features. The figure is shown for iteration $j=1$
@@ 5 Experiments
We first provide results on our newly formulated visual classifiers. We use the ImageNet-1K dataset due to the widespread availability of visual classifiers trained and evaluated on it. We apply our method on a diverse set of 40 visual classifiers. For CNNs, we consider the following family of models (each with several variants): Residual Networks (ResNets) (He et al., 2015), Wide ResNets (Zagoruyko & Komodakis, 2016), ResNeXts (Xie et al., 2016), ShuffleNetv2 (Ma et al., 2018), EfficientNetv2 (Tan & Le, 2021), Densely Connected Networks (DenseNets) (Huang et al., 2016), ConvNeXts (Liu et al., 2022) and ConvNeXtv2 (Woo et al., 2023). For Transformers, we consider the following family of models (each with several variants): Vision Transformers (ViTs) (Dosovitskiy et al., 2021), DINOv2 (Oquab et al., 2024), BeiT (Bao et al., 2022), the hybrid Convolution-Vision Transformer CvT (Wu et al., 2021), Swin Transformer (Liu et al., 2021b) and Swin Transformer v2 (Liu et al., 2021a). All models are pretrained on ImageNet-1K from the PyTorch [^3] https://pytorch.org/vision/stable/models.html and HuggingFace (Wolf et al., 2020) libraries. Models with the subscript pt indicate that the model was pretrained on ImageNet-21k before being finetuned on ImageNet-1K. Models with a subscript v2 are trained following the updated PyTorch training recipe (Vryniotis, 2021). Finally, BEiT, DINOv2 and ConvNeXtv2 are pretrained in a self-supervised manner before being finetuned on ImageNet-1k. Both the pretrained classifier and text encoder remain frozen, only the MLP is trained on the ImageNet training set following Eq. 1.
Performance is evaluated using the same protocol and dataset splits as the original classifier, specifically the 50,000 validation split of ImageNet-1K. For the text encoder, we use the MiniLM Sentence Encoder (Wang et al., 2020), with ablation studies on alternative text encoders shown in Section B of the Appendix. For implementation details, we refer to Section D of the Appendix. Results are presented in Table 1. We report the replicated Top-1 accuracy of the re-formulated classifier with our method in the first column, the original Top-1 accuracy of the classifier in the second column, and the difference between them ($\Delta$) in the last column. As can be seen, the loss in performance as indicated by $\Delta$ is minimal, with an average drop in performance of approximately $0.2$ points across all models. We refer to Appendix Section C for results on other models.
Model Top-1 Orig. $\Delta$
ResNet50 75.80 76.13 --0.33
@@
DenseNet161 77.04 77.14 --0.10
DenseNet169 75.46 75.60 --0.14
EfficientNetv2-S 84.04 84.23 --0.19
EfficientNetv2-M 84.95 85.11 --0.16
[…]
ConvNeXt-Small 83.42 83.62 --0.20
ConvNeXt-Base 83.88 84.06 --0.18
[…]
ResNeXt50-32x4d 77.44 77.62 --0.18
[…]
ResNeXt101-64x4d 83.13 83.25 --0.12
ResNeXt101-32x8d 79.10 79.31 --0.21
ViT-B/16 80.70 81.07 --0.37
[…]
ViT-L/32 76.72 76.97 --0.25
ViT-L/16 79.56 79.66 --0.10
Swin-Base 83.22 83.58 --0.36
Swinv2-Base 83.72 84.11 --0.39
BeiT-B/16 84.54 85.06 --0.52
BeiT-L/16 87.22 87.34 --0.12
DINOv2-B 84.40 84.22 +0.18
ConvNeXtV2-B 84.56 84.73 --0.17
@@
Table 1: Performance comparison of our re-formulated classifiers for several models. Top-1 indicates our results of the new formulation, and Orig. denotes the original Top-1 accuracy. $\Delta$ represents the difference between them $\Delta$ = Top-1 - Orig.
Zero-Shot Concept Bottleneck Models: We report CBM evaluation results on the ImageNet validation set using the top-1 accuracy in Table 2. As baselines, we compare with LF-CBMs (Oikarinen et al., 2023), LaBo (Yang et al., 2022), CDM (Panousis et al., 2023) and the recent state-of-the-art DN-CBMs (Rao et al., 2024). All these methods are supervised and use CLIP-based models. Our zero-shot CBMs (ZS-CBMs) outperforms all the supervised CBMs and sets a new state-of-the-art. Notably, even a simple ResNet-50 classifier trained solely on ImageNet already outperforms the CBM for the significantly more powerful ResNet-50 CLIP model trained on 400M samples. The best results are obtained by the ConvNeXtv2 and ViT-L/16 models, which achieve a top-1 accuracy of $86.3$. All models show close to original accuracy, which means we can transform any classifier to be inherently interpretable without much loss in performance.
Method Model Top-1
Supervised CBMs
LF-CBM CLIP ResNet50 67.5
LF-CBM CLIP ViT-B/16 75.4
LaBo CLIP ResNet50 68.9
LaBo CLIP ViT-B/16 78.9
CDM CLIP ResNet50 72.2
CDM CLIP ViT-B/16 79.3
DCLIP CLIP ViT-B/16 68.0
DN-CBM CLIP ResNet50 72.9
DN-CBM CLIP ViT-B/16 79.5
Zero-Shot CBMs (Ours)
ZS-CBM ResNet50 73.9
@@
ZS-CBM ResNet101 75.3
[…]
ZS-CBM WideResNet50 76.9
@@
ZS-CBM DenseNet121 69.9
ZS-CBM DenseNet161 75.2
ZS-CBM EfficientNetv2-S 83.0
ZS-CBM EfficientNetv2-M 83.9
ZS-CBM ConvNeXt-Small 81.9
ZS-CBM ConvNeXt-Base 82.8
ZS-CBM ViT-B/32 73.3
ZS-CBM ViT-B/16 79.3
[…]
ZS-CBM Swin-Base 82.2
ZS-CBM Swinv2-Base 82.6
[…]
ZS-CBM BeiT-B/16 83.0
ZS-CBM DINOv2-B 82.6
@@
Table 2: Performance of Zero-Shot CBMs (ZS-CBM) on ImageNet validation set for several classifiers.
Figure 5: Qualitative examples of our zero-shot CBMs. We show the top-detected concepts, each with their corresponding importance score to the prediction. All examples use the 20K concept set, except for the first which uses the ImageNet LF-CBM concept set
Figure 6: Global class-wise interpretability analysis with our Concept Bottleneck Model. We highlight in yellow the top concepts in ”hammerhead shark” that are not present in ”tiger shark”, and therefore distinctive to ”hammerhead shark”.
In Figure 5, we present qualitative examples of a selection from the top concepts responsible for the prediction, along with their weight importance. The weight importance is calculated by multiplying the concept activation with its corresponding weight to the predicted class. By observing the second example, we see that the image is predicted as a “goose” because it has duck-like features, it is in the size of a swan, and it has bird and pigeon-like features. Similarly, we can see that the third image is a “spotted salamander” due to its lizard-like structure, a head that looks like the frog’s head, a snake-like tail and spots like a leopard. In the last image, we can see that the prediction is “ostrich” due to its camel-like hump, mammal-like size, piegeon-like body, the horse-riding-like feature, and is a type of a bird. In Figure 6, we present a probability distribution of global class-wise concepts. These are concepts detected for all images of a specific class, along with their frequency. We consider two semantically similar classes but distinctively different: “hammerhead shark” and a “tiger shark”. We highlight in yellow the top concepts in “hammerhead shark” that are not present in “tiger shark”. These concepts are “harpoon” and “lobster hammer”, both which are distinctive to the head of the hammerhead shark and drive its prediction.
Figure 7: Qualitative examples of decoded visual features from different visual classifiers, into natural text.
@@
Zero-Shot Decoding of Visual Features: We now evaluate the performance of the textually decoded visual features. Since the ground-truth content of visual features is unknown, we rely on COCO captions (Lin et al., 2014) for evaluation. While these captions may not perfectly correspond to the actual information represented by a visual feature vector—potentially differing from what a human annotator expects—they provide a reasonable approximation for this purpose. Note that the COCO dataset differs in distribution than ImageNet, as a single image may contain many objects, interactions between them, and potentially categories not included in ImageNet (e.g., person). Therefore, it also serves as a way to evaluate generalization of our method to other datasets, given that we only used the ImageNet class names for training. Since we do not train any model on the ground-truth image captions provided by COCO, we use zero-shot image captioning as a benchmark. We present results on the widely used “Karpathy test split” with various vision classifiers.
@@
As baselines, we compare our approach against existing methods in zero-shot image captioning, specifically ZeroCap (Tewel et al., 2021) and ConZIC (Zeng et al., 2023), both which use CLIP. For evaluation, we employ standard natural language generation (NLG) metrics: BLEU-4 (B@4) (Papineni et al., 2002), METEOR (M) (Banerjee & Lavie, 2005), ROUGE-L (R-L) (Lin, 2004), CIDEr (C) (Vedantam et al., 2014), and SPICE (S) (Anderson et al., 2016). Results are shown Table 3. ConvNeXtv2 achieves state-of-the-art performance on CIDEr and SPICE, the two most critical metrics for evaluating image captioning systems. Even with a simple ResNet-50 vision encoder trained on ImageNet-1K (1.2 million images), our approach outperforms the baseline methods on CIDEr and SPICE, despite the latter utilizing the significantly more powerful CLIP vision encoder, trained on 400 million image-text pairs. We also present qualitative examples of the decoded visual features for different models in Figure 7. We can see from the first example, that BeiT-L/16 contains both the vegetables and dog in its features, while ConvNexTv2 only focuses on the vegetables, and ViT-B/16 and DINOv2 only focus on on the dog. Although some generations lacks semantic correctness (e.g., who was born in a dog), they are still meaningful enough for humans to understand and reason. This application allows a user-friendly way of interpreting visual features, applicable even to a layman user.
@@
Table 3: Zero-Shot Image Captioning Performance
@@
Table 4: Composition Captioning Performance across models.
@@
Note that our results in Table 3 are outperformed by the baseline ZeroCap on the BLEU-4 (B4) and METEOR (M) metrics. However, it is important to note that B4 and M are n-gram overlap-based metrics. They assume that the generated caption follows a specific structure and style. We verify this by revisiting an old image captioning paradigm termed as compositional captioning (Kulkarni et al., 2011), and later revived with deep learning methods (Lu et al., 2018). In compositional captioning, a set of image-grounded concepts (such as attributes, objects and verbs) are first detected, and a language model is then used to compose them into a natural sounding sentence. With the current advancements of Large Language Models (LLMs), we can use an LLM as a compositioner. Specifically, we detect the top concepts and verbs to the image using the concept discovery method introduced in Section 4.1 and shown in Figure 3, and feed them, along with their similarity scores, to an LLM. For the concepts, we use the same concept set as in Section 4.1. We also add a list of the most common verbs [^4] https://github.com/datmt/English-Words-Updated in English to the pool. This allows us to cover all possible words and interactions. We prompt the LLM to utilize the provided information to compose a sentence, given a very few in-context examples from the COCO captioning training set (in our experiments, we use 9 examples). This allows us to generate sentences adhering to a specific style and structure. For this experiment, we used GPT4o-mini (OpenAI, 2024) as our LLM, as it is fast and cost-efficient. In the prompt, we explicitly instruct the LLM to refrain from reasoning or generating content based on its own knowledge or assumptions, and that all its outputs must be strictly grounded in the provided concepts, verbs, and score importances. We provide details on the exact prompt we used in Appendix Section G. Results are shown in Table 4. While results on CiDEr and SPICE are incremental compared to results in Table 3, the n-gram metrics (B4, M and R-L) are boosted, which verifies our hypothesis about the low scores of B4 and M in Table 3 compared to baseline methods. Also note that since the LLM is generic and can adapt to any style and structure, compositional captioning is especially useful for generating sentences in a specific style (e.g., humorous, positive, or negative).
6 Conclusion and Future Directions

### v2

@@ 4.2 Zero-Shot Decoding of Visual Features into Text
In this application, we aim to decode the visual feature vector $f$ for an image $I$, given by $f=F_{v}(I)$, into a natural language sentence. This offers a text-based interpretation of what the visual features contain. We adapt the method introduced in [3] for our purpose. Specifically, we first project the visual feature vector $f$ using the MLP to obtain $\tilde{f}$. That is, $\tilde{f}=\text{MLP}(f)$. Since $\tilde{f}$ is now in the same space as the text encoder $T$, we can measure its association to any encoded text. We utilize an off-the-shelf pretrained language decoder model (e.g., GPT-2), denoted as $G$, to generate open-ended text. We keep $G$ frozen to maintain its language generation capabilities and instead use prefix-tuning [20] to guide $G$ to generate a text that maximizes the similarity with the transformed visual feature vector $\tilde{f}$. Specifically, we attach a set of learnable “virtual" tokens to $G$. Denote the generated output text of $G$ for one iteration as $h^{j}$, where $j$ represents the iteration number. $h^{j}$ is encoded with the text encoder $T$, to yield a vector $y^{j}$, That is, $y^{j}=T(h^{j})$. As $y^{j}$ and $\tilde{f}$ are now in the same embedding space, we maximize the cosine similarity between them in order to update the learnable tokens. We perform this process for several iterations. An overview of this process is shown in Figure 3(b). As this process is not the core contribution of our work, we leave more details to Section H of the appendix, which also includes a detailed illustration in Figure 2.
@@ 5 Experiments
We first provide results on our newly formulated visual classifiers. We use the challenging ImageNet-1K dataset due to the widespread publicly available visual classifiers trained and evaluated on it. We apply our method on a diverse set of 40 visual classifiers. For CNNs, we consider the following family of models (each with several variants): Residual Networks (ResNets) [21], Wide ResNets [22], ResNeXts [23], ShuffleNetv2 [24], EfficientNetv2 [25], Densely Connected Networks (DenseNets) [26], ConvNeXts [27] and ConvNeXtv2 [28]. For Transformers, we consider the following family of models (each with several variants): Vision Transformers (ViTs) [29], DINOv2 [30], BeiT [31], the hybrid Convolution-Vision Transformer CvT [32], Swin Transformer [33] and Swin Transformer v2 [34]. All models are pretrained on ImageNet-1K from the PyTorch [35] and HuggingFace [36] libraries. Models with the subscript pt indicate that the model was pretrained on ImageNet-21k before being finetuned on ImageNet-1K. Models with a subscript v2 are trained following the updated PyTorch training recipe [37]. Finally, BEiT, DINOv2 and ConvNeXtv2 are pretrained in a self-supervised manner before being finetuned on ImageNet-1k. Both the pretrained classifier and text encoder remain frozen, only the MLP is trained on the ImageNet training set following Eq. 1.
Performance is evaluated using the same protocol and dataset splits as the original classifier, specifically the 50,000 validation split of ImageNet-1K. For the text encoder, we use the MiniLM Sentence Encoder [38], with ablation studies on other text encoders presented in Section D of the Appendix. Results are presented in Table 1. We report the replicated Top-1 accuracy of the re-formulated classifier with our method in the first column, the original Top-1 accuracy of the classifier in the second column, and the difference between them ($\Delta$) in the last column. As can be seen, the loss in performance as indicated by $\Delta$ is minimal, with an average drop in performance of approximately $0.2$ points across all models. We also perform ablation studies on the MLP to verify its effectiveness in Section C of the Appendix. We refer to Appendix Section F for results on other models and to Section G for implementation details.
Model Top-1 Orig. $\Delta$ Model Top-1 Orig. $\Delta$
ResNet50 75.80 76.13 --0.33 ResNeXt50-32x4dv2 80.79 80.88 --0.09
@@
DenseNet169 75.46 75.60 --0.14 Swin-Base 83.22 83.58 --0.36
EfficientNetv2-S 84.04 84.23 --0.19 Swinv2-Base 83.72 84.11 --0.39
EfficientNetv2-M 84.95 85.11 --0.16 BeiT-B/16 84.54 85.06 --0.52
@@
ConvNeXt-Small 83.42 83.62 --0.20 DINOv2-B 84.40 84.22 +0.18
ConvNeXt-Base 83.88 84.06 --0.18 ConvNeXtV2-B 84.56 84.73 --0.17
@@
Table 1: Performance comparison of our re-formulated classifiers for several models. Top-1 indicates our results of the new formulation, and Orig. denotes the original Top-1 accuracy. $\Delta$ represents the difference between them ($\Delta$ = Top-1 -- Orig).
Zero-Shot Concept Bottleneck Models: We report CBM evaluation results on the ImageNet validation set using the top-1 accuracy in Table 2. As baselines, we compare with 9 methods: LF-CBMs [15], LaBo [16], CDM [17], DCLIP [39], DN-CBMs [18], DCBM [40], Multi-Modal (MM) CBMs [41], CF-CBM [42] and SALF-CBM [43]. All these methods are supervised and use CLIP-based models either for computing the concept activations or as a backbone. Our zero-shot CBMs (ZS-CBMs) outperforms all the supervised CBMs and sets a new state-of-the-art. Notably, even a simple ResNet-50 classifier trained solely on ImageNet already outperforms the CBM for the significantly more powerful ResNet-50 CLIP model trained on 400M samples. The best results are obtained by the ConvNeXtv2 model, which achieve a top-1 accuracy of $86.4$. All models show close to original accuracy, which means we can transform any classifier to be inherently interpretable without much loss in performance.
Figure 4: Qualitative examples of our zero-shot CBMs. We show the top-detected concepts, each with their corresponding importance score to the prediction.
Effectivenss of Concept Interventions: As a supplementary experiment, we also report concept intervention results on our ZS-CBMs in Section E of the Appendix, showing the effectiveness of the concepts, how we can mitigate biases, debug models and fix their reasoning by explicitly intervening in the concepts of the bottleneck layer to control predictions.
Supervised CBMs
Method Model Top-1 Method Model Top-1
LF-CBM CLIP ResNet50 67.5 DN-CBM CLIP ViT-B/16 79.5
LF-CBM CLIP ViT-B/16 75.4 SALF-CBM-S CLIP ViT-B/16 75.3
LaBo CLIP ResNet50 68.9 SALF-CBM-D CLIP ViT-B/16 76.3
LaBo CLIP ViT-B/16 78.9 DCBM-SAM2 CLIP ViT-L/14 77.9
CDM CLIP ResNet50 72.2 DCBM-GDINO CLIP ViT-L/14 77.4
CDM CLIP ViT-B/16 79.3 DCBM-RCNN CLIP ViT-L/14 77.8
DCLIP CLIP ViT-B/32 63.0 MM-CBM CLIP ViT-B/16 77.9
DCLIP CLIP ViT-B/16 68.0 MM-ProtoSim CLIP ViT-B/16 78.8
DCLIP CLIP ViT-L/14 75.0 CF-CBM-H CLIP ViT-B/16 77.4
DN-CBM CLIP ResNet50 72.9 CF-CBM-L CLIP ViT-B/16 78.4
Zero-Shot CBMs (Ours)
Method Model Top-1 Method Model Top-1
ZS-CBM ResNet50 73.9 ZS-CBM ViT-B/32 73.3
@@
Table 2: Performance of Supervised and Zero-Shot CBMs on the ImageNet validation set.
In Figure 4, we present qualitative examples of a selection from the top concepts responsible for the prediction, along with their weight importance. The weight importance is calculated by multiplying the concept activation with its corresponding weight to the predicted class. We use various concept sets to demonstrate the flexibility of our method to any desired concept set directly at test time (on-the-fly), as this process simply involves encoding the chosen concept set using the text encoder. All examples use the LF-CBM concept set [15]. By observing the first example, the image is predicted as a “scorpion" because it is has a lizard-like overall body, it is a venomous, desert animal with claws on its legs, and has large claws on its hand that look like a crab. Interestingly, in the second example, the top-detected concept is a “lift arm on the side.” Although this is not the primary feature defining a dumbbell, it reflects a well-documented bias in the literature [44] regarding the “dumbbell” class. Because most training images for this class show a dumbbell being lifted by an arm, the classifier not only learns to recognize the dumbbell but also associates it with the hand or arm that lifts it. With our method, we can obtain a textual interpretation of the biases that the original classifier learns. We also provide qualitative examples of global class-wise concepts detected in Section K of the appendix.
@@
Zero-Shot Decoding of Visual Features: We now evaluate the performance of the textually decoded visual features. Since the ground-truth content of visual features is unknown—that is, we do not know exactly what the model uses for reasoning—we rely on COCO captions [45] for evaluation. While these captions may not perfectly correspond to the actual information represented by a visual feature vector—potentially differing from what a human annotator expects—they provide a reasonable approximation for this purpose, and represent the best available alternative for evaluation. Note that the COCO dataset differs in distribution than ImageNet, as a single image may contain many objects, interactions between them, and potentially categories not included in ImageNet (e.g., person). Therefore, it also serves as a way to evaluate generalization of our method to other datasets, given that we only used the ImageNet class names for training. Since we do not train any model on the ground-truth image captions provided by COCO, we use zero-shot image captioning as a benchmark.
@@
We present results on the widely used “Karpathy test split" with various vision classifiers. As baselines, we compare our approach against existing methods in zero-shot image captioning, specifically ZeroCap [3] and ConZIC [46], both which use CLIP. For evaluation, we employ standard natural language generation metrics: BLEU-4 (B@4) [47], METEOR (M) [48], ROUGE-L (R-L) [49], CIDEr (C) [50], and SPICE (S) [51]. Results are shown in Table 3(a). ConvNeXtv2 achieves state-of-the-art performance on CIDEr and SPICE, the two most critical metrics for evaluating image captioning systems. Even with a simple ResNet-50 vision encoder trained on ImageNet-1K (1.2 million images), our approach outperforms the baseline methods on CIDEr and SPICE, despite the latter utilizing the significantly more powerful CLIP vision encoder, trained on 400 million image-text pairs. We also present qualitative examples of the decoded visual features for different models in Figure 5. From the first example, BeiT-L/16 captures features of both the vegetables and the dog, whereas ConvNeXtV2 captures only the vegetables. In contrast, ViT-B/16 and DINOv2 focus exclusively on the dog and its characteristics. Although some generations lacks semantic correctness (e.g., who was born in a dog), they are still meaningful enough for humans to understand and reason. This application allows human-friendly interpretations of visual features, understandable even to a layman user.
@@
(a) Zero-Shot Image Captioning Performance
@@
(b) Zero-Shot Compositional Captioning Performance
Table 3: Comparison of Zero-Shot Captioning Performance
Figure 5: Qualitative examples of decoded visual features from different visual classifiers, into text.
@@
Note that our results in Table 3(a) are outperformed by the baseline ZeroCap on the BLEU-4 (B4) and METEOR (M) metrics. However, it is important to note that B4 and M are n-gram overlap-based metrics. They assume that the generated caption follows a specific structure and style. We verify this by revisiting an old image captioning paradigm termed as compositional captioning [52], and later revived with deep learning methods [53]. In compositional captioning, a set of image-grounded concepts (such as attributes, objects and verbs) are first detected, and a language model is then used to compose them into a natural sounding sentence. With the current advancements of Large Language Models (LLMs), we can use an LLM as a compositioner. Specifically, we detect the top concepts and verbs to the image using the concept discovery method introduced in Section 4.1 and shown in Figure 3(a), and feed them, along with their similarity scores, to an LLM. For the concepts, we use the same concept set as in Section 4.1. We also add a list of the most common verbs [54] in English to the pool. This allows us to cover all possible words and interactions. We prompt the LLM to utilize the provided information to compose a sentence, given a very few in-context examples from the COCO captioning training set (in our experiments, we use 9 examples). This allows us to generate sentences adhering to a specific style and structure. For this experiment, we used GPT4o-mini [55] as our LLM, as it is fast and cost-efficient. In the prompt, we explicitly instruct the LLM to refrain from reasoning or generating content based on its own knowledge or assumptions, and that all its outputs must be strictly grounded in the provided concepts, verbs, and score importances. We provide details on the exact prompt we used in Appendix Section J. Results are shown in Table 3(b). While results on CiDEr and SPICE are incremental compared to results in Table 3(a), the n-gram metrics (B4, M and R-L) are boosted, which verifies our hypothesis about the low scores of B4 and M in Table 3(a) compared to baseline methods. Also note that, since the LLM is generic and can adapt to any style and structure, compositional captioning is especially useful for generating captions tailored to specific domains (see Section J.1 in the appendix).
6 Conclusion